OpenStudy (anonymous):

Solve. x2 – x – 6 = 0 A. {3, –2} B. {6, –1} C. {2, –3} D. {–2, –3}

OpenStudy (jadeishere):

Is there a certain method you have to use?

OpenStudy (anonymous):

no

OpenStudy (jadeishere):

Okay I'll do the factoring way so x^2 means it'll be (x ) (x) -x = -1x and that is the sum of two numbers, and -6 is the product, of those two numbers. 2 and 3 works in this way, -3 + 2 = -1 so you get -x for that and -3 * 2 = -6 so (x +2)(x - 3) so (2, -3)

OpenStudy (anonymous):

Solve. 2x2 + x – 1 = 2

OpenStudy (jadeishere):

a = 2x2 b = 1x c = -3 if you do the quadratic formula, it'll be easier in this one. for me, anyway

OpenStudy (jadeishere):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (jadeishere):

so plug in the given values of a, b and c to the formula and you get \[x = \ \frac{ -1 \pm \sqrt{1^2 - 4(2)(-3)}}{ 4 }\]

OpenStudy (the_fizicx99):

Jadey I believe they're referring to complete the square >.> it's faster <,<

OpenStudy (jadeishere):

I know two ways that are faster for me >.<

OpenStudy (the_fizicx99):

:o

OpenStudy (jadeishere):

I already know the answer I was just going to work it out for milbes

OpenStudy (the_fizicx99):

Mental Math ♦o♦

OpenStudy (jadeishere):

Lol it's pretty easy. >.<

OpenStudy (the_fizicx99):

If you know it by heart it is <,<

OpenStudy (jadeishere):

\[x = \frac{ -1 \pm \sqrt{25} }{ 4}\]

OpenStudy (jadeishere):

Lol

OpenStudy (anonymous):

thats what i got then i divided it

OpenStudy (jadeishere):

Welp... I probably should have done it differently then, it doesn't make sense if it's -1 >.< Boy oh boy i'm out of it, aren't I?

OpenStudy (anonymous):

gtg can u explain it fast

OpenStudy (anonymous):

@Xmoses1

OpenStudy (jadeishere):

It's B

OpenStudy (jadeishere):

Sorry but just trust me on this one, it's definitely B, -1 - 5 = -6 -6 / 4 = -3/2